A hyperbola eq. is 9y2-4x2=36

Find the value of e

Grade:11

3 Answers

Akash Kumar Dutta
11 years ago

Dear Anurag,

The question you have posted is exremely easy.

a^2=4,b^2=9
a^2(e^2 - 1)=b^2.
putting the value we get. e=(13)^1/2
                                              ____
                                               2
(ANS).

Please post some tough questions next time.

Regards.

anurag kushwaha
11 years ago

Ans is √13/3

Akash Kumar Dutta
11 years ago

Dear Anurag,

I think i have made a mistake.
can you please display the solution.

Regards.